OpenStudy (anonymous):

5. Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. -x – 8y = 43 4x + 32y = 172

OpenStudy (anonymous):

-x-8y=43(1) X+8y=43 4x+32y=172 5x+40y=215 9x+72y=215

OpenStudy (jlastino):

Hint: Multiply the first equation by 4 and let's see what happens when you add the two equations

OpenStudy (jlastino):

Dude when you solve by elimination method, you multiply a number that, when you add/subtract it to another equation, would eliminate a variable from the given equations